total assets 3,540,000. total equity 1222000. liabilities current. current ratio .65. how much are xurrent assets.
Answer #1

Sol: We are given following information :-
Current ratio = 0.65 : 1

Total Assets = $ 3,540,000

Total Equity = $ 1,222,000

Assuming that on the Liabilities side of Balance Sheet, there are no other Liabilities than Current Liabilities.

So Current Liabilities = Total Assets - Total Equity

Current Liabilities = $ 3,540,000 - $ 1,222,000 = $ 2,318,000

Now we have Current Ratio = (Current Assets/ Current Liabilities)

We have the values of Current ratio = 0.65
Current Liabilities = $ 2,318,000
& let Current assets = x

So Current ratio = 0.65 = ( x / 2,318,000)

So x = 2,318,000 * 0.65 = $ 1,506,700.

So we have, Current Assets = $ 1506700